Work on all problems. Copy and paste the excel into the answer.4. (20 pts) Expected values versus most likely values. The radial part of the wave function for the hydrogen atom in the 2p s
Answer #1

(a)R_{2p}(r)=Are^{-\frac{r}{2a_B}}

Normalising the wavefunction

\int_0^\infty r^2 R^2_{2p}(r)dr=1

A^2\int_0^\infty r^4 e^{-\frac{r}{a_B}}dr=1

Let, \frac{r}{a_B}=x , dr=a_B dx

A^2 a_B^5\int_0^\infty x^4 e^{-x}dx=1

A^2 a_B^5*4!=1

A=\sqrt \frac{1}{4!a_B^5}

Hence normalised wavefunction is

R_{2p}(r)= \sqrt \frac{1}{4!a_B^5}re^{-\frac{r}{2a_B}}

(b)

The graph of the function is:

te

The maxima of the function will lie at r=2a_B . So, it the most likely radius isr=2a_B

(c) R_{2p}(r)= \sqrt \frac{1}{4!a_B^5}re^{-\frac{r}{2a_B}}

For most likely radius,

\frac{\mathrm{d} R_{2p}}{\mathrm{d} r}=0

\frac{\mathrm{d} }{\mathrm{d} r} \left (\sqrt \frac{1}{4!a_B^5}re^{-\frac{r}{2a_B}} \right )=0

\sqrt \frac{1}{4!a_B^5} \frac{\mathrm{d} }{\mathrm{d} r} \left (re^{-\frac{r}{2a_B}} \right )=0

\frac{\mathrm{d} }{\mathrm{d} r} \left (re^{-\frac{r}{2a_B}} \right )=0

e^{-\frac{r}{2a_B}}-\frac{r}{2a_B}e^{-\frac{r}{2a_B}}=0

\left (1-\frac{r}{2a_B} \right )e^{-\frac{r}{2a_B}}=0

\left (1-\frac{r}{2a_B} \right ) =0

Hence, r=2a_B is the most likely radius.

(d)

Average radius is given by:

<r>=\int_0^\infty r*r^2 R^2_{2p}(r)dr

<r>=\int_0^\infty r^5 \frac{1}{4!a_B^5}e^{-\frac{r}{a_B}}dr

<r>= \frac{1}{4!a_B^5}\int_0^\infty r^5 e^{-\frac{r}{a_B}}dr

Let, \frac{r}{a_B}=x , dr=a_B dx

<r>= \frac{a_B^6}{4!a_B^5}\int_0^\infty x^5 e^{-x}dx

<r>= \frac{a_B^6}{4!a_B^5}5!

<r>= 5a_B